OpenStudy (anonymous):

can someone help me with 2y^2+9y+9

Directrix (directrix):

What do you want to do with 2y^2+9y+9 ? What are the instructions?

OpenStudy (anonymous):

factor the trinomial

Directrix (directrix):

What are two integers that multiply to 18 and add to 9? We need those.

OpenStudy (anonymous):

why would it be 18 just wondering

Directrix (directrix):

In the trinomial, ax^2 + bx + c, where a is not zero, to begin to factor by grouping, find two numbers that multiply to a and c such that those numbers also add to b. In your problem, a = 2 and c = 9. So, we look for two numbers that multiply to ac and add to b.

Directrix (directrix):

If you want to guess and hope to get the factors by trial and error, you might get lucky. I can't help with that.

OpenStudy (anonymous):

Ok thank you well my homework will have to wait I will be on later

OpenStudy (anonymous):

ok sorry about that im back so 18 and 9 is 6 and 3

Directrix (directrix):

2y² + 9y + 9 = 2y² + 3y + 6y + 9 = y ( 2y + 3) + 3(2y + 3) = [Note that there are common factors of (2y + 3)] (2y + 3) (y + 3) @jen89

OpenStudy (anonymous):

ok thank you

Directrix (directrix):

You're welcome.

OpenStudy (anonymous):

can you help me with another one please

OpenStudy (anonymous):

it is 4c^2-9c+5

Directrix (directrix):

What are two integers that multiply to 20 and also add to -9 ?

OpenStudy (anonymous):

5 and -4

OpenStudy (anonymous):

so do i just add those into the problem

Directrix (directrix):

5 and -4 = 1 Try again. The two numbers must add to -9

OpenStudy (anonymous):

4 and 5

Directrix (directrix):

4 + 5 = 9. The two numbers we want must add to -9. You are close. Try again.

OpenStudy (anonymous):

im stuck on this one i get -8 or -1 not -9

Directrix (directrix):

Let's try -4 and -5. 4c² - 9c + 5 = 4c² -5c - 4c + 5 = c(4c - 5) - (4c - 5) = c(4c - 5) - 1(4c - 5) = (4c - 5) (c - 1)

OpenStudy (anonymous):

ok i get it now